# RadioActive: 3D Radiological Interactive Segmentation Benchmark

### v2

@@ 7 Additional Results
Table 4: Experimental results simulating a realistic clinician’s effort. ’center PPV’ stands for Point Per Volume, that was sampled from the center of the target object. Sampling from the center performs better than sampling a random point. SegVol trained on CT images from the D2 HanSeg dataset, while we tested on MRI images from the same patient cohort.
Iterations Prompter Model Interactions D1 D2 D3 D4 D5 D6 D7 D8 D9 D10 Average
0 1PPS + 1PPS Refine SAM 1X/1X 0.81 2.66 3.13 2.06 1.42 1.0 7.86 12.47 3.42 3.33 3.82
1 1PPS + 1PPS Refine SAM 1X/1X 1.02 2.94 4.46 3.22 1.97 1.13 8.73 13.81 3.5 3.56 4.43
2 1PPS + 1PPS Refine SAM 1X/1X 1.11 2.97 4.92 3.66 2.5 1.41 10.0 15.55 3.77 3.82 4.97
3 1PPS + 1PPS Refine SAM 1X/1X 1.18 2.37 5.24 4.41 3.03 1.55 11.06 16.97 4.02 4.0 5.38
4 1PPS + 1PPS Refine SAM 1X/1X 1.18 2.03 4.64 4.95 3.33 1.62 11.88 17.7 4.16 4.38 5.59
5 1PPS + 1PPS Refine SAM 1X/1X 1.11 2.2 3.3 5.56 3.37 1.43 12.57 18.25 4.09 4.32 5.62
0 1PPS + 1PPS Refine SAM2 1X/1X 1.25 2.74 4.08 3.9 3.07 1.51 9.6 16.35 3.97 3.83 5.03
1 1PPS + 1PPS Refine SAM2 1X/1X 2.96 4.99 8.69 6.2 5.64 3.32 17.94 24.31 5.49 5.78 8.53
2 1PPS + 1PPS Refine SAM2 1X/1X 5.25 6.83 11.68 8.03 7.64 5.04 24.56 29.52 7.39 7.83 11.38
3 1PPS + 1PPS Refine SAM2 1X/1X 7.53 8.8 15.02 10.01 9.52 6.87 29.11 33.53 9.06 9.59 13.9
4 1PPS + 1PPS Refine SAM2 1X/1X 9.81 10.51 18.31 11.99 11.09 8.24 33.22 48.49 10.45 11.54 17.37
5 1PPS + 1PPS Refine SAM2 1X/1X 12.14 12.15 21.33 13.52 12.56 9.21 37.04 51.38 11.72 13.5 19.45
0 1PPS + 1PPS Refine SamMed 2D 1X/1X 10.72 26.3 24.55 28.83 22.76 9.71 30.83 31.24 26.05 20.31 23.13
1 1PPS + 1PPS Refine SamMed 2D 1X/1X 12.44 29.97 28.54 31.35 24.3 11.78 39.25 37.54 31.41 21.45 26.8
2 1PPS + 1PPS Refine SamMed 2D 1X/1X 12.49 31.19 30.25 32.28 24.65 12.41 43.74 40.46 33.49 22.07 28.3
3 1PPS + 1PPS Refine SamMed 2D 1X/1X 12.59 31.93 31.49 32.85 24.98 12.88 46.69 42.32 34.71 22.6 29.31
4 1PPS + 1PPS Refine SamMed 2D 1X/1X 12.78 32.51 32.59 33.31 25.33 13.31 48.82 43.68 35.59 23.05 30.1
5 1PPS + 1PPS Refine SamMed 2D 1X/1X 12.93 32.98 33.67 33.74 25.76 13.67 50.46 44.81 36.29 23.49 30.78
0 1PPS + 1PPS Refine * SAM 1X/1X 0.81 2.66 3.13 2.06 1.42 1.0 7.86 12.47 3.42 3.33 3.82
1 1PPS + 1PPS Refine * SAM 1X/1X 0.06 0.27 0.63 0.6 0.15 0.08 3.34 3.38 1.45 0.61 1.06
2 1PPS + 1PPS Refine * SAM 1X/1X 0.33 0.95 1.55 1.64 0.76 0.67 6.2 11.11 2.61 2.45 2.83
3 1PPS + 1PPS Refine * SAM 1X/1X 0.22 1.06 1.56 1.42 0.86 0.51 5.16 7.74 2.14 1.54 2.22
4 1PPS + 1PPS Refine * SAM 1X/1X 0.38 1.08 1.73 1.4 0.5 0.57 6.07 10.85 2.61 2.3 2.75
5 1PPS + 1PPS Refine * SAM 1X/1X 0.2 0.93 1.34 1.29 0.62 0.58 5.53 8.34 2.07 1.69 2.26
0 1PPS + 1PPS Refine * SAM2 1X/1X 1.25 2.74 4.08 3.9 3.07 1.51 9.6 16.35 3.97 3.83 5.03
1 1PPS + 1PPS Refine * SAM2 1X/1X 1.33 2.94 5.86 4.34 3.44 1.76 11.73 18.45 4.13 3.74 5.77
2 1PPS + 1PPS Refine * SAM2 1X/1X 1.17 3.02 6.24 4.62 3.53 1.64 11.58 17.89 4.3 3.66 5.77
3 1PPS + 1PPS Refine * SAM2 1X/1X 1.29 3.05 6.31 4.93 3.69 1.73 11.78 17.81 4.34 3.67 5.86
4 1PPS + 1PPS Refine * SAM2 1X/1X 1.31 3.22 6.57 4.97 3.79 1.75 11.82 17.5 4.37 3.84 5.91
5 1PPS + 1PPS Refine * SAM2 1X/1X 1.23 3.21 6.6 4.86 3.74 1.76 11.97 17.78 4.4 3.8 5.93
0 1PPS + 1PPS Refine * SamMed 2D 1X/1X 10.72 26.3 24.55 28.83 22.76 9.71 30.83 31.24 26.05 20.31 23.13
1 1PPS + 1PPS Refine * SamMed 2D 1X/1X 10.45 28.62 27.4 31.37 23.26 10.47 38.52 37.22 30.63 20.04 25.8
2 1PPS + 1PPS Refine * SamMed 2D 1X/1X 9.88 28.71 28.05 31.7 22.67 10.37 41.79 38.9 31.55 19.47 26.31
3 1PPS + 1PPS Refine * SamMed 2D 1X/1X 9.6 28.45 28.26 31.42 22.18 10.58 43.5 40.16 31.9 18.97 26.5
4 1PPS + 1PPS Refine * SamMed 2D 1X/1X 9.2 28.17 28.29 31.08 21.78 10.44 44.71 40.81 31.99 18.52 26.5
5 1PPS + 1PPS Refine * SamMed 2D 1X/1X 8.99 27.84 28.26 30.56 21.49 10.54 45.45 41.43 31.99 18.2 26.48
0 1PPS + Scribble Refine SAM 1/3 0.81 2.66 3.13 2.06 1.42 1.0 7.86 12.47 3.42 3.33 3.82
1 1PPS + Scribble Refine SAM 1/3 0.89 2.65 4.53 3.24 2.11 1.18 7.48 11.65 3.55 3.68 4.1
2 1PPS + Scribble Refine SAM 1/3 0.93 2.76 4.45 3.54 2.35 1.37 8.01 11.24 3.52 3.41 4.16
3 1PPS + Scribble Refine SAM 1/3 0.96 2.12 3.5 3.49 2.09 1.4 8.01 10.6 3.4 3.15 3.87
4 1PPS + Scribble Refine SAM 1/3 0.92 1.94 2.08 3.43 1.94 1.4 8.06 9.84 3.27 3.01 3.59
5 1PPS + Scribble Refine SAM 1/3 0.88 1.69 1.4 3.43 1.87 1.14 7.66 9.54 3.14 2.56 3.33
0 1PPS + Scribble Refine SAM2 1/3 1.25 2.74 4.08 3.9 3.07 1.51 9.6 16.35 3.97 3.83 5.03
1 1PPS + Scribble Refine SAM2 1/3 1.74 3.88 8.13 5.39 5.97 3.01 16.14 18.28 4.66 5.07 7.23
2 1PPS + Scribble Refine SAM2 1/3 2.18 4.42 10.05 5.84 7.57 4.14 19.28 19.15 4.86 5.75 8.32
3 1PPS + Scribble Refine SAM2 1/3 2.52 4.8 10.93 6.16 8.2 5.06 21.87 25.97 4.9 6.17 9.66
4 1PPS + Scribble Refine SAM2 1/3 2.83 5.03 12.37 6.26 8.75 5.88 24.46 27.84 4.86 6.45 10.47
5 1PPS + Scribble Refine SAM2 1/3 3.1 5.3 13.38 6.26 9.08 6.93 25.81 30.02 4.86 6.85 11.16
0 1PPS + Scribble Refine SamMed 2D 1/3 10.72 26.3 24.55 28.83 22.76 9.71 30.83 31.24 26.05 20.31 23.13
1 1PPS + Scribble Refine SamMed 2D 1/3 11.8 28.84 26.82 29.72 23.07 11.32 37.22 36.7 29.87 20.7 25.61
2 1PPS + Scribble Refine SamMed 2D 1/3 11.71 29.29 27.88 29.5 22.7 11.63 40.64 38.77 31.61 20.65 26.44
3 1PPS + Scribble Refine SamMed 2D 1/3 11.55 29.52 28.26 29.15 22.67 11.68 42.57 39.92 32.62 20.68 26.86
4 1PPS + Scribble Refine SamMed 2D 1/3 11.49 29.62 28.33 29.19 22.6 11.76 43.94 40.71 33.25 20.74 27.16
5 1PPS + Scribble Refine SamMed 2D 1/3 11.46 29.75 28.72 28.87 22.52 11.88 44.93 41.32 33.68 20.83 27.4
0 3B Inter + Scribble Refine* SAM 6/3 13.13 54.04 63.34 58.64 64.26 71.35 65.52 61.78 25.58 59.23 53.69
1 3B Inter + Scribble Refine* SAM 6/3 0.99 3.66 3.24 3.53 2.33 6.14 7.78 11.88 4.68 6.43 5.07
2 3B Inter + Scribble Refine* SAM 6/3 0.85 1.95 1.28 3.17 9.04 5.28 7.49 14.4 4.14 7.43 5.51
3 3B Inter + Scribble Refine* SAM 6/3 0.8 4.11 2.95 2.19 3.13 3.36 7.14 9.85 4.02 4.67 4.22
4 3B Inter + Scribble Refine* SAM 6/3 0.71 2.45 2.01 2.21 2.28 1.89 6.08 9.09 3.22 3.57 3.35
5 3B Inter + Scribble Refine* SAM 6/3 0.76 1.99 2.46 2.51 2.16 1.42 6.15 9.34 3.19 3.47 3.34
0 3B Inter + Scribble Refine* SAM2 6/3 69.78 57.59 67.71 69.08 69.06 73.11 68.35 63.24 37.21 63.27 63.84
1 3B Inter + Scribble Refine* SAM2 6/3 10.03 20.17 25.37 19.13 23.65 18.12 25.37 27.66 11.34 15.93 19.68
2 3B Inter + Scribble Refine* SAM2 6/3 5.07 6.98 10.53 7.31 8.35 3.99 20.7 19.58 5.61 6.41 9.45
3 3B Inter + Scribble Refine* SAM2 6/3 4.23 4.43 7.8 6.93 5.64 3.72 14.43 16.35 4.89 5.34 7.38
4 3B Inter + Scribble Refine* SAM2 6/3 3.99 3.84 6.96 7.22 4.86 3.53 13.37 15.02 4.56 4.9 6.83
5 3B Inter + Scribble Refine* SAM2 6/3 3.41 3.7 6.75 5.7 3.99 2.79 12.49 13.59 4.3 4.47 6.12
0 3B Inter + Scribble Refine* SamMed 2D 6/3 27.83 42.22 52.7 60.66 44.76 38.63 55.31 55.81 33.46 38.66 45.0
1 3B Inter + Scribble Refine* SamMed 2D 6/3 28.4 44.72 57.53 63.33 45.28 41.42 61.61 59.73 40.44 38.8 48.13
2 3B Inter + Scribble Refine* SamMed 2D 6/3 24.29 42.41 53.02 59.24 40.27 35.69 62.86 58.85 41.23 34.93 45.28
3 3B Inter + Scribble Refine* SamMed 2D 6/3 20.91 39.69 47.48 52.73 35.55 28.2 62.08 54.89 40.02 31.59 41.31
4 3B Inter + Scribble Refine* SamMed 2D 6/3 18.76 37.29 41.46 46.5 31.73 23.98 58.46 51.86 38.07 28.74 37.69
5 3B Inter + Scribble Refine* SamMed 2D 6/3 17.07 35.06 37.39 41.62 29.1 20.8 53.9 49.27 36.27 26.66 34.71
0 3P Inter + Scribble Refine SAM 5/3 0.84 3.6 3.5 2.17 1.7 1.02 8.3 12.61 3.64 4.5 4.19
1 3P Inter + Scribble Refine SAM 5/3 0.89 3.29 4.59 3.58 2.41 1.21 7.96 11.68 3.73 4.39 4.37
2 3P Inter + Scribble Refine SAM 5/3 0.98 3.19 5.18 4.21 2.65 1.44 8.64 11.54 3.68 4.24 4.58
3 3P Inter + Scribble Refine SAM 5/3 1.03 2.5 3.61 4.14 2.52 1.5 9.12 10.95 3.64 3.71 4.27
4 3P Inter + Scribble Refine SAM 5/3 1.03 2.04 2.23 4.14 2.31 1.48 8.91 10.4 3.48 3.32 3.93
5 3P Inter + Scribble Refine SAM 5/3 0.99 1.89 1.5 4.17 1.87 1.32 8.16 9.75 3.33 3.02 3.6
0 3P Inter + Scribble Refine SAM2 5/3 1.38 3.68 4.73 4.1 3.37 1.63 10.44 16.98 4.4 5.13 5.58
1 3P Inter + Scribble Refine SAM2 5/3 1.87 5.18 8.77 6.04 6.8 3.47 16.57 19.63 5.09 6.31 7.97
2 3P Inter + Scribble Refine SAM2 5/3 2.37 5.83 11.33 7.01 9.24 4.62 22.39 21.17 5.41 7.22 9.66
3 3P Inter + Scribble Refine SAM2 5/3 2.84 6.27 13.65 7.62 10.71 5.9 23.75 27.73 5.48 7.65 11.16
4 3P Inter + Scribble Refine SAM2 5/3 3.33 6.52 14.69 7.83 11.49 6.96 27.72 28.86 5.52 8.05 12.1
5 3P Inter + Scribble Refine SAM2 5/3 3.5 6.7 15.79 7.97 12.18 7.65 30.01 30.82 5.56 8.47 12.86
0 3P Inter + Scribble Refine SamMed 2D 5/3 11.71 29.54 28.46 34.42 26.17 10.88 36.81 35.65 28.27 23.31 26.52
1 3P Inter + Scribble Refine SamMed 2D 5/3 12.72 31.26 30.27 34.05 26.02 12.06 40.92 38.13 31.14 23.12 27.97
2 3P Inter + Scribble Refine SamMed 2D 5/3 12.46 31.66 30.58 33.13 25.57 12.19 42.79 39.25 32.55 22.88 28.31
3 3P Inter + Scribble Refine SamMed 2D 5/3 12.25 31.79 30.53 32.57 25.24 12.28 43.89 39.98 33.35 22.73 28.46
4 3P Inter + Scribble Refine SamMed 2D 5/3 12.13 31.83 30.42 32.12 25.05 12.33 44.89 40.52 33.84 22.63 28.58
5 3P Inter + Scribble Refine SamMed 2D 5/3 12.06 31.86 30.42 31.66 24.9 12.41 45.85 40.97 34.22 22.61 28.7
Table 5: Interactive refinement results for 2D models across 5 iterations. The initial prediction is made either using a single point per slice or one of our proposed prompting schemes. If the previous point prompt is not reused during refinement, indicated by a *, the performance drops. This also explains the drop in performance for the 3 Box interpolation after the initial prompt. The unrealistic slice-wise refinement (1 interaction per slice) is only slightly better than our proposed scribble refinement method (3 interactions).
Iterations Prompter Model Interactions D1 D2 D3 D4 D5 D6 D7 D8 D9 D10 Average
0 1 center PPV + 1 PPV Refine* SamMed 3D 1/1 2.03 12.15 24.11 26.9 15.09 19.64 72.66 53.03 50.4 26.51 30.25
1 1 center PPV + 1 PPV Refine* SamMed 3D 1/1 5.89 14.07 33.03 35.67 14.42 24.93 64.15 72.35 42.64 24.86 33.2
2 1 center PPV + 1 PPV Refine* SamMed 3D 1/1 5.67 13.49 29.52 28.92 14.34 21.4 64.29 62.44 42.03 25.16 30.73
3 1 center PPV + 1 PPV Refine* SamMed 3D 1/1 5.57 13.64 31.57 35.51 14.21 22.82 63.35 65.94 41.12 24.79 31.85
4 1 center PPV + 1 PPV Refine* SamMed 3D 1/1 5.81 13.51 31.44 31.08 14.11 21.38 65.28 63.38 40.68 24.82 31.15
5 1 center PPV + 1 PPV Refine* SamMed 3D 1/1 5.46 13.56 32.19 33.49 14.18 24.28 62.12 65.2 40.41 25.12 31.6
[…]

### v3

@@ 8.8 ScribblePromt
Table 4: Experimental results simulating a realistic clinician’s effort. ’center PPV’ refers to points sampled at equal axial distances along the five-point interpolation prompter line, which approximates the centerline of the target volume. Sampling from the center performs better than sampling a random point only for the first point. SegVol trained on CT images from the D2 HanSeg dataset, while we tested on MRI images from the same patient cohort.
Model Promter Interactions Iteration D1 D2 D3 D4 D5 D6 D7 D8 D9 D10 Average
SAM 1PPS + 1PPS Refine 1X/1X 0 1.89 3.36 3.06 4.21 1.43 0.92 11.06 15.4 28.52 4.89 7.47
SAM 1PPS + 1PPS Refine 1X/1X 1 0.1 1.11 0.17 1.07 0.01 0.02 0.43 0.71 5.57 0.75 0.99
SAM 1PPS + 1PPS Refine 1X/1X 2 0.45 2.0 2.08 2.31 1.01 0.63 6.64 11.62 13.0 3.26 4.3
SAM 1PPS + 1PPS Refine 1X/1X 3 0.63 2.8 1.13 2.46 0.88 0.42 4.93 6.76 9.58 2.18 3.18
SAM 1PPS + 1PPS Refine 1X/1X 4 0.43 2.08 1.95 1.76 0.67 0.6 6.3 11.17 12.4 3.18 4.05
SAM 1PPS + 1PPS Refine 1X/1X 5 0.4 2.21 1.47 1.81 0.65 0.53 5.08 7.84 8.52 2.31 3.08
SAM2 1PPS + 1PPS Refine 1X/1X 0 3.36 4.99 4.41 5.91 3.77 1.62 15.64 22.36 45.86 6.79 11.47
SAM2 1PPS + 1PPS Refine 1X/1X 1 3.32 5.79 6.72 6.83 4.95 1.99 23.76 28.73 47.73 6.98 13.68
SAM2 1PPS + 1PPS Refine 1X/1X 2 2.41 6.3 7.63 8.2 5.23 1.99 25.8 29.07 48.78 7.36 14.28
SAM2 1PPS + 1PPS Refine 1X/1X 3 2.4 7.11 8.9 8.39 5.35 1.98 28.33 30.34 50.66 7.59 15.1
SAM2 1PPS + 1PPS Refine 1X/1X 4 2.52 7.15 9.36 8.15 5.45 1.85 29.92 30.69 51.29 7.52 15.39
SAM2 1PPS + 1PPS Refine 1X/1X 5 3.51 7.43 9.39 7.78 5.89 1.91 30.41 31.13 52.16 7.62 15.72
SamMed 2D 1PPS + 1PPS Refine 1X/1X 0 18.28 33.98 44.01 41.44 35.97 18.79 59.09 49.33 62.66 24.28 38.78
SamMed 2D 1PPS + 1PPS Refine 1X/1X 1 17.56 36.81 44.97 45.81 36.19 20.8 69.07 56.35 67.4 23.5 41.85
SamMed 2D 1PPS + 1PPS Refine 1X/1X 2 16.99 37.43 44.89 46.33 35.01 21.54 73.2 59.28 68.65 22.7 42.6
SamMed 2D 1PPS + 1PPS Refine 1X/1X 3 16.57 37.5 44.52 46.4 34.15 21.76 75.4 60.96 69.23 22.01 42.85
SamMed 2D 1PPS + 1PPS Refine 1X/1X 4 16.05 37.41 44.22 46.02 33.28 21.9 76.63 62.01 69.57 21.43 42.85
SamMed 2D 1PPS + 1PPS Refine 1X/1X 5 15.78 37.2 43.95 45.55 32.82 21.93 77.34 62.85 69.78 21.03 42.82
ScribblePrompt 1PPS + 1PPS Refine 1X/1X 0 25.29 32.9 47.38 49.09 26.43 4.37 49.73 45.4 55.47 35.12 37.12
ScribblePrompt 1PPS + 1PPS Refine 1X/1X 1 4.85 14.18 22.53 20.7 1.79 0.47 20.28 13.97 32.68 16.44 14.79
ScribblePrompt 1PPS + 1PPS Refine 1X/1X 2 11.69 21.84 28.48 25.57 16.35 2.98 34.11 31.91 46.36 26.93 24.62
ScribblePrompt 1PPS + 1PPS Refine 1X/1X 3 6.4 16.48 23.47 23.54 6.1 0.72 22.08 17.91 33.56 19.55 16.98
ScribblePrompt 1PPS + 1PPS Refine 1X/1X 4 11.75 21.51 28.54 25.55 15.73 2.99 33.37 31.02 46.18 27.24 24.39
ScribblePrompt 1PPS + 1PPS Refine 1X/1X 5 6.68 16.52 23.32 22.21 6.69 0.91 20.91 17.86 32.23 17.91 16.52
SAM 1PPS + 1PPS Refine* 1X/1X 0 1.89 3.36 3.06 4.21 1.43 0.92 11.06 15.4 28.52 4.89 7.47
SAM 1PPS + 1PPS Refine* 1X/1X 1 2.31 3.95 4.8 4.31 2.93 1.27 14.36 18.66 17.76 5.55 7.59
SAM 1PPS + 1PPS Refine* 1X/1X 2 2.96 5.1 5.88 6.83 3.63 1.68 17.71 22.64 25.27 6.72 9.84
SAM 1PPS + 1PPS Refine* 1X/1X 3 3.67 4.76 6.7 8.48 3.97 1.88 21.2 24.34 29.31 7.19 11.15
SAM 1PPS + 1PPS Refine* 1X/1X 4 3.84 4.27 5.99 10.49 4.77 1.85 22.79 25.3 32.39 7.33 11.9
SAM 1PPS + 1PPS Refine* 1X/1X 5 2.8 4.09 4.64 12.52 5.79 1.88 24.89 25.93 32.2 7.09 12.18
SAM2 1PPS + 1PPS Refine* 1X/1X 0 3.36 4.99 4.41 5.91 3.77 1.62 15.64 22.36 45.86 6.79 11.47
SAM2 1PPS + 1PPS Refine* 1X/1X 1 5.86 13.4 13.61 13.8 9.17 4.58 43.18 42.1 60.63 12.8 21.91
SAM2 1PPS + 1PPS Refine* 1X/1X 2 10.11 20.08 19.99 20.62 14.69 8.49 59.67 51.88 70.53 18.52 29.46
SAM2 1PPS + 1PPS Refine* 1X/1X 3 14.5 25.49 25.19 26.3 19.94 11.6 66.94 58.3 75.98 23.44 34.77
SAM2 1PPS + 1PPS Refine* 1X/1X 4 19.3 30.35 28.84 30.21 25.06 14.15 72.21 87.37 79.46 27.32 41.43
SAM2 1PPS + 1PPS Refine* 1X/1X 5 23.71 34.35 32.83 33.5 29.61 15.81 74.65 90.05 81.96 30.76 44.72
SamMed 2D 1PPS + 1PPS Refine* 1X/1X 0 18.28 33.98 44.01 41.44 35.97 18.79 59.09 49.33 62.66 24.28 38.78
SamMed 2D 1PPS + 1PPS Refine* 1X/1X 1 18.65 37.46 44.63 45.02 37.47 20.36 68.45 56.05 67.19 25.01 42.03
SamMed 2D 1PPS + 1PPS Refine* 1X/1X 2 18.93 39.03 45.11 46.16 37.44 20.91 72.76 59.35 69.13 25.53 43.44
SamMed 2D 1PPS + 1PPS Refine* 1X/1X 3 19.25 40.03 45.77 46.89 37.48 21.31 75.23 61.37 70.4 26.03 44.37
SamMed 2D 1PPS + 1PPS Refine* 1X/1X 4 19.6 40.85 46.52 47.45 37.68 21.69 76.85 62.81 71.34 26.5 45.13
SamMed 2D 1PPS + 1PPS Refine* 1X/1X 5 19.89 41.53 47.25 47.84 37.96 22.03 77.88 63.9 72.08 26.93 45.73
ScribblePrompt 1PPS + 1PPS Refine* 1X/1X 0 25.29 32.9 47.38 49.09 26.43 4.37 49.73 45.4 55.47 35.12 37.12
ScribblePrompt 1PPS + 1PPS Refine* 1X/1X 1 28.79 36.93 55.37 56.62 30.29 7.0 51.12 38.99 64.36 41.32 41.08
ScribblePrompt 1PPS + 1PPS Refine* 1X/1X 2 35.33 42.85 62.06 66.91 39.29 11.94 59.68 44.7 69.95 45.81 47.85
ScribblePrompt 1PPS + 1PPS Refine* 1X/1X 3 39.56 47.12 65.91 71.03 43.54 18.35 64.93 50.5 72.84 48.29 52.21
ScribblePrompt 1PPS + 1PPS Refine* 1X/1X 4 41.75 49.85 67.77 74.24 46.29 26.52 70.31 55.53 74.36 49.92 55.65
ScribblePrompt 1PPS + 1PPS Refine* 1X/1X 5 43.34 51.55 68.72 75.34 48.11 34.68 73.27 58.74 75.14 50.79 57.97
SAM 1PPS + Scribble Refine 1x/3 0 1.89 3.36 3.06 4.21 1.43 0.92 11.06 15.4 28.52 4.89 7.47
SAM 1PPS + Scribble Refine 1x/3 1 0.38 4.5 1.17 8.51 1.21 0.24 9.78 9.87 24.8 6.21 6.67
SAM 1PPS + Scribble Refine 1x/3 2 0.7 3.82 2.41 12.91 1.69 0.76 9.02 10.73 26.84 5.25 7.41
SAM 1PPS + Scribble Refine 1x/3 3 0.59 4.09 1.63 6.96 1.33 0.56 8.61 10.79 27.88 6.28 6.87
SAM 1PPS + Scribble Refine 1x/3 4 0.68 3.49 1.43 8.82 0.9 0.61 9.57 10.02 25.32 5.42 6.63
SAM 1PPS + Scribble Refine 1x/3 5 0.53 3.41 1.51 7.99 1.41 0.55 7.75 10.07 26.61 6.69 6.65
SAM2 1PPS + Scribble Refine 1x/3 0 3.36 4.99 4.41 5.91 3.77 1.62 15.64 22.36 45.86 6.79 11.47
SAM2 1PPS + Scribble Refine 1x/3 1 3.64 5.89 5.5 6.05 3.84 2.12 17.01 21.49 45.67 7.88 11.91
SAM2 1PPS + Scribble Refine 1x/3 2 3.65 6.07 5.7 5.63 2.91 2.14 17.8 20.17 43.8 7.78 11.57
SAM2 1PPS + Scribble Refine 1x/3 3 3.45 5.89 5.27 6.32 2.71 2.22 18.64 18.56 41.95 7.64 11.27
SAM2 1PPS + Scribble Refine 1x/3 4 3.4 5.57 4.88 6.02 2.58 2.13 19.27 17.91 41.23 7.5 11.05
SAM2 1PPS + Scribble Refine 1x/3 5 3.38 5.41 4.94 5.52 2.34 1.77 18.69 18.46 40.69 7.45 10.86
SamMed 2D 1PPS + Scribble Refine 1x/3 0 18.28 33.98 44.01 41.44 35.97 18.79 59.09 49.33 62.66 24.28 38.78
SamMed 2D 1PPS + Scribble Refine 1x/3 1 17.38 35.9 44.18 44.43 34.82 18.99 67.64 55.93 66.38 23.62 40.93
SamMed 2D 1PPS + Scribble Refine 1x/3 2 16.42 35.5 43.49 43.96 32.82 19.22 72.01 57.86 66.72 22.85 41.08
SamMed 2D 1PPS + Scribble Refine 1x/3 3 16.04 35.15 42.47 43.96 31.55 18.67 71.25 58.44 66.48 22.0 40.6
SamMed 2D 1PPS + Scribble Refine 1x/3 4 15.38 34.65 42.06 43.88 29.83 18.88 71.7 58.67 66.18 21.39 40.26
SamMed 2D 1PPS + Scribble Refine 1x/3 5 15.04 34.22 41.44 43.88 28.87 19.59 71.46 59.17 65.89 20.74 40.03
ScribblePrompt 1PPS + Scribble Refine 1x/3 0 25.29 32.9 47.37 49.09 26.43 4.37 49.73 45.4 55.47 35.12 37.12
ScribblePrompt 1PPS + Scribble Refine 1x/3 1 6.3 22.19 38.63 39.72 10.24 1.16 36.22 32.16 52.76 24.5 26.39
ScribblePrompt 1PPS + Scribble Refine 1x/3 2 13.26 24.48 38.53 33.5 13.26 2.88 36.67 31.76 51.3 32.37 27.8
ScribblePrompt 1PPS + Scribble Refine 1x/3 3 7.99 21.89 35.94 36.36 10.49 1.21 34.76 27.56 49.24 26.1 25.15
ScribblePrompt 1PPS + Scribble Refine 1x/3 4 11.13 23.92 39.96 34.97 12.58 2.39 34.69 29.82 50.13 31.82 27.14
ScribblePrompt 1PPS + Scribble Refine 1x/3 5 7.99 20.98 36.75 32.74 9.9 1.16 33.85 31.63 48.47 27.55 25.1
SAM 1PPS + Scribble Refine* 1X/3 0 1.89 3.36 3.06 4.21 1.43 0.92 11.06 15.4 28.52 4.89 7.47
SAM 1PPS + Scribble Refine* 1X/3 1 2.64 3.6 4.63 5.15 2.14 1.1 10.25 13.86 37.1 6.2 8.67
SAM 1PPS + Scribble Refine* 1X/3 2 2.84 4.49 5.27 8.53 2.63 1.45 11.35 15.05 39.68 6.5 9.78
SAM 1PPS + Scribble Refine* 1X/3 3 2.97 4.3 4.74 10.05 2.48 1.63 12.45 15.15 38.94 6.1 9.88
SAM 1PPS + Scribble Refine* 1X/3 4 3.06 3.43 3.42 12.12 2.22 1.64 13.03 14.33 35.61 5.28 9.41
SAM 1PPS + Scribble Refine* 1X/3 5 2.15 2.96 2.43 11.85 2.52 1.52 13.4 14.67 33.81 4.66 9.0
SAM2 1PPS + Scribble Refine* 1X/3 0 3.36 4.99 4.41 5.91 3.77 1.62 15.64 22.36 45.86 6.79 11.47
SAM2 1PPS + Scribble Refine* 1X/3 1 4.88 10.1 10.19 14.09 9.61 4.0 34.23 33.66 56.19 11.36 18.83
SAM2 1PPS + Scribble Refine* 1X/3 2 5.7 12.98 14.7 20.06 13.52 6.56 41.64 41.38 59.46 13.97 23.0
SAM2 1PPS + Scribble Refine* 1X/3 3 6.86 14.98 18.73 21.55 15.91 8.75 49.62 46.04 61.29 15.48 25.92
SAM2 1PPS + Scribble Refine* 1X/3 4 7.43 16.58 22.06 23.03 17.47 10.33 53.22 50.35 62.16 16.6 27.92
SAM2 1PPS + Scribble Refine* 1X/3 5 7.86 17.84 25.14 23.8 19.11 11.97 55.59 52.72 62.95 17.77 29.47
SamMed 2D 1PPS + Scribble Refine* 1X/3 0 18.28 33.98 44.01 41.44 35.97 18.79 59.09 49.33 62.66 24.28 38.78
SamMed 2D 1PPS + Scribble Refine* 1X/3 1 18.19 36.17 43.27 43.15 35.71 18.97 66.82 54.92 66.13 24.23 40.76
SamMed 2D 1PPS + Scribble Refine* 1X/3 2 18.09 37.12 42.97 42.87 35.33 18.99 69.63 57.03 67.39 24.0 41.34
SamMed 2D 1PPS + Scribble Refine* 1X/3 3 18.03 37.6 42.73 42.56 35.15 19.0 71.31 58.23 68.06 23.9 41.66
SamMed 2D 1PPS + Scribble Refine* 1X/3 4 18.05 37.94 42.53 42.25 34.99 19.07 72.41 59.14 68.56 23.8 41.87
SamMed 2D 1PPS + Scribble Refine* 1X/3 5 18.08 38.18 42.48 42.05 34.88 19.08 73.18 59.9 68.93 23.72 42.05
ScribblePrompt 1PPS + Scribble Refine* 1X/3 0 25.29 32.9 47.37 49.09 26.43 4.37 49.73 45.4 55.47 35.12 37.12
ScribblePrompt 1PPS + Scribble Refine* 1X/3 1 27.54 33.07 50.97 42.56 24.39 12.87 42.75 36.19 59.75 38.81 36.89
ScribblePrompt 1PPS + Scribble Refine* 1X/3 2 32.12 35.91 52.98 50.54 27.9 20.19 45.32 37.63 61.23 41.94 40.58
ScribblePrompt 1PPS + Scribble Refine* 1X/3 3 34.98 39.06 55.9 57.26 31.82 28.71 51.78 41.82 62.79 43.86 44.8
ScribblePrompt 1PPS + Scribble Refine* 1X/3 4 36.65 41.64 57.53 62.96 35.13 34.67 57.44 45.05 64.59 45.99 48.17
ScribblePrompt 1PPS + Scribble Refine* 1X/3 5 37.76 43.84 59.31 67.05 37.72 41.1 60.27 47.49 65.83 47.42 50.78
SAM 3B Inter + Scribble Refine* 6/3 0 73.31 60.53 80.06 NaN 75.75 77.85 86.54 76.3 62.48 67.48 73.37
SAM 3B Inter + Scribble Refine* 6/3 1 NaN 57.97 76.5 NaN 72.43 75.4 82.98 72.22 63.08 69.85 71.3
SAM 3B Inter + Scribble Refine* 6/3 2 NaN 56.94 74.58 NaN 72.64 75.02 80.92 72.78 65.59 71.25 71.21
SAM 3B Inter + Scribble Refine* 6/3 3 NaN 56.17 73.97 NaN 73.29 74.03 81.56 72.53 67.25 71.56 71.3
SAM 3B Inter + Scribble Refine* 6/3 4 NaN 55.13 71.9 NaN 73.41 72.74 82.01 70.69 68.39 71.67 70.74
SAM 3B Inter + Scribble Refine* 6/3 5 NaN 53.44 69.1 NaN 72.93 70.31 82.46 69.29 69.31 70.93 69.72
SAM2 3B Inter + Scribble Refine* 6/3 0 77.29 65.06 80.2 82.62 78.6 80.43 86.24 78.41 58.8 71.4 75.9
SAM2 3B Inter + Scribble Refine* 6/3 1 77.9 66.22 82.03 84.28 79.08 81.23 86.87 80.31 61.82 71.5 77.12
SAM2 3B Inter + Scribble Refine* 6/3 2 79.22 67.37 82.98 85.49 80.69 83.6 87.5 81.39 63.31 73.8 78.53
[…]